# Kazhdan groups of dimension $16$ with an uncountable range of second $\ell^{2}$-Betti numbers

### v1

@@ 6 Proof of the main theorem
    There is a finite set of finite subgroups $\mathcal{E}_{i}$ of $G_{i/m}$, such that for every unitary $G_{i/m}$-representation $V$, the second homology is
    $H_{2}(G_{i/m};V)=\bigoplus_{E\in\mathcal{E}_{i}}V_{E}.$
Then the final group $G_{m/m}$ will be the group $G_{1}$ in the statement of the lemma, with $\mathcal{E}=\mathcal{E}_{m}$ and $x_{j}=x_{j}^{m}$. The base case for $i=0$ is $G$ itself, with $\mathcal{E}=\emptyset$, where we do not need to check anything for the first three items, and Item 4 follows from Proposition 3.8 and Theorem 3.10.
Suppose by induction that $G_{i/m}$ has been constructed. If $x_{i+1}^{i}$ is torsion, then it is $q$-torsion, so we set $G_{(i+1)/m}=G_{i/m}$, $x^{i+1}_{j}=x^{i}_{j}$ for all $j$, and $\mathcal{E}_{i+1}=\mathcal{E}_{i}$. Otherwise, consider the infinite virtually cyclic group $E_{G_{i/m}}(x_{i+1})$. This has a unique maximal finite normal subgroup of order a power of $q$, and the quotient is isomorphic to $\mathbb{Z}$ (or possibly to $D_{\infty}$, in case $q=2$). We replace $x_{i+1}^{i}$ with a root $x^{i+1}_{i+1}$ that generates the $\mathbb{Z}$ quotient (or the index-$2$ subgroup of the $D_{\infty}$ quotient), so that $E_{G_{i/m}}(x^{i+1}_{i+1}):\langle x^{i+1}_{i+1}\rangle]$ is a power of $q$. We now let $G_{(i+1)/m}$ be the group-theoretic Dehn filling for the triple $(G_{i/m},E_{G_{i/m}}(x^{i+1}_{i+1}),\langle(x^{i+1}_{i+1})^{k}\rangle)$, where $k$ is a sufficiently large power of $q$. By Proposition 4.7, we see that $G_{(i+1)/m}$ satisfies Items 1 and 3, choosing $x^{i+1}_{j}=x^{i}_{j}$ for all $j\neq i+1$. We also choose $k$ to be large enough so that the orders of $x_{1}^{i},\ldots,x_{i}^{i}$ and their images $x_{1}^{i+1},\ldots,x_{i}^{i+1}$ have the same order, and each finite subgroup from $\mathcal{E}_{i}$ embeds into $G_{(i+1)/m}$. Moreover, by Proposition 4.7 again, this triple is Cohen–Lyndon, so Corollary 4.3 applies, and proves Item 2. Finally, let $E\coloneqq E_{G_{i/m}}(x_{i+1}^{i+1})/\langle(x_{i+1}^{i+1})^{k}\rangle$. Corollary 4.3 and property $(T)$ for $G_{i/m}$ give a short exact sequence
$0\to H_{2}(G_{i/m};V)\to H_{2}(G_{(i+1)/m};V)\to V_{E}\to 0,$
so setting $\mathcal{E}_{i+1}=\mathcal{E}_{i}\cup\{E\}$, we obtain Item 4 and conclude. ∎
@@
We proceed as in the proof of Theorem 6.1, using the prime $p\neq q$, but starting with the group $G_{1}$ from Lemma 6.2. The result will be a group $G_{\alpha}$ with the same properties as in Theorem 6.1, with the exception that cohomological dimension now has to be taken modulo $\{p,q\}$, and $b^{(2)}_{2}(G_{\alpha})$ is computed starting from
$b^{(2)}_{2}(G_{1}\to G_{\alpha})=\dim_{\mathcal{N}(G_{\alpha})}\Bigl(\bigoplus_{E\in\mathcal{E}}\mathcal{N}(G_{\alpha})_{E}\Bigr)=\sum\limits_{E\in\mathcal{E}}|E|^{-1}.$
Since this starting value is independent of $\alpha$, the same computations go through. For the finiteness of the $\ell^{2}$-Betti numbers of $G_{\alpha}$ we refer to Remark 6.3.
In each induction step in the proof of Theorem 6.1, we chose arbitrary special elements $g_{i}\in G_{i}$. We will now explain how to use the full power of Proposition 4.8 in the construction so that this time the colimit is simple. Enumerate the set $\mathcal{X}\coloneqq\{x_{1},\ldots,x_{m}\}\times G_{1}$. When the group $G_{i}$ has been built, let $(x,h)$ denote the $i$-th element of $\mathcal{X}$. If $h$ is trivial, then we pick $g_{i}$ to be an arbitrary special element, and proceed as before. Otherwise, we use Proposition 4.8 to pick $g_{i}\in x^{-1}\left<\!\left<h\right>\!\right>_{G_{i}}$. We claim that now $x\in\left<\!\left<h\right>\!\right>_{G_{i+1}}\leqslant G_{i+1}$. Indeed, consider the image of $x$ in the quotient $G_{i+1}/\left<\!\left<h\right>\!\right>_{G_{i+1}}$. On the one hand, it has order a power of $q$, because $x$ has order a power of $q$ in $G_{1}$. On the other hand, it equals the image of $g_{i}$, which has order a power of $p$ in $G_{i+1}$, by construction of the quotient $G_{i}\to G_{i+1}$. Because $p\neq q$, we see that $x$ must be trivial in $G_{i+1}/\left<\!\left<h\right>\!\right>_{G_{i+1}}$, which proves our claim.
As a result, for every $(x,h)\in\mathcal{X}$, either $h$ is trivial in $G_{\alpha}$, or $x$ lies in the normal closure of $h$ in $G_{\alpha}$. Because $x$ ranges over a generating set for $G_{1}$ and $h$ ranges over $G_{1}$, this shows that every non-trivial element of $G_{\alpha}$ normally generates, that is, $G_{\alpha}$ is simple. ∎
@@
Note that the constructions give some information about higher $\ell^{2}$-Betti numbers as well. Indeed, $G_{\alpha}$ is obtained from a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, denoted $G$, by a sequence of group-theoretic Dehn fillings satisfying Corollary 4.3. Using this and Proposition 2.2 we have
$b^{(2)}_{n}(G_{\alpha})=b^{(2)}_{n}(G\to G_{\alpha})$
for all $n\geqslant 3$. Since $G$ is of type $F$, this implies that all $\ell^{2}$-Betti numbers of $G_{\alpha}$ are finite. We know from Theorem 3.10 that $G$ has property $(T_{3})$, so Corollary 3.9 gives $b^{(2)}_{3}(G_{\alpha})=0$.
Moreover, $G$ satisfies the Singer conjecture [Lüc02, Corollary 5.16 on p. 231], that is $b^{(2)}_{n}(G)=0$ for all $n\neq 8$ and $b^{(2)}_{8}(G)=\chi\geqslant 1$. This does not give information about $b^{(2)}_{n}(G\to G_{\alpha})$, however, assuming Lück’s approximation conjecture [Lüc02, Chapter 13], and repeating the construction ensuring that $G\to G_{\alpha}$ is injective on a large ball centred at the identity, we would have that $b^{(2)}_{n}(G\to G_{\alpha})$ is close to $b^{(2)}_{n}(G)$. It would then follow that $b^{(2)}_{n}(G_{\alpha})$ is close to 0, for $n\geqslant 4,n\neq 8$, and close to $\chi$ for $n=8$.
@@
Taking $V=\mathbb{C}$ to be the trivial unitary representation, the proofs of Theorems 6.1 and A, and Proposition 2.2, show that $H_{2}(G_{\alpha};\mathbb{C})$ is infinite-dimensional. It follows that the groups $G_{\alpha}$ do not have type $FP_{2}(\mathbb{Q})$.
@@ 7 An elementary construction
We conclude with the elementary construction giving D. We start with two basic product formulas.
@@
Taking $\dim_{G_{x}}$ and $n\leqslant 3$, the first and last term give $2b^{(2)}_{n}(R_{x}\times F_{2}^{3})=2b^{(2)}_{n-1}(R_{x}\times F_{2}^{3})=0$, by Proposition 3.11, so
@@
Let $(G_{x})_{x\in\mathbb{R}_{>0}}$ be as in D. Then the groups $G_{x}\ast\mathbb{Z}$ are pairwise non-measure equivalent.

### v2

@@ 6 Reals and simple groups
    There is a finite set of finite subgroups $\mathcal{E}_{i}$ of $G^{i}$, such that for every unitary $G^{i}$-representation $V$, the second homology is
    $H_{2}(G^{i};V)=\bigoplus_{E\in\mathcal{E}_{i}}V_{E},$
    and moreover
    $\sum\limits_{E\in\mathcal{E}_{i}}|E|^{-1}<\varepsilon.$
Then the final group $G^{m}$ will be the group $G_{0}$ in the statement of the lemma, with $\mathcal{E}=\mathcal{E}_{m}$ and $a_{j}=a_{j}^{m}$. The base case for $i=0$ is $G$ itself, with $\mathcal{E}=\emptyset$, where we do not need to check anything for the first three items, and Item 4 follows from Proposition 3.8 and Theorem 3.10.
Suppose by induction that $G^{i}$ has been constructed. If $a_{i+1}^{i}$ is torsion, then it is $q$-torsion, so we set $G^{i+1}=G^{i}$, $a^{i+1}_{j}=a^{i}_{j}$ for all $j$, and $\mathcal{E}_{i+1}=\mathcal{E}_{i}$. Otherwise, consider the infinite virtually cyclic group $E_{G^{i}}(a_{i+1})$. This has a unique maximal finite normal subgroup of order a power of $q$, and the quotient is isomorphic to $\mathbb{Z}$ (or possibly to $D_{\infty}$, in case $q=2$). We replace $a_{i+1}^{i}$ with a root $a^{i+1}_{i+1}$ that generates the $\mathbb{Z}$ quotient (or the index-$2$ subgroup of the $D_{\infty}$ quotient), so that $E_{G^{i}}(a^{i+1}_{i+1}):\langle a^{i+1}_{i+1}\rangle]$ is a power of $q$. We now let $G^{i+1}$ be the group-theoretic Dehn filling for the triple $(G^{i},E_{G^{i}}(a^{i+1}_{i+1}),\langle(a^{i+1}_{i+1})^{k}\rangle)$, where $k$ is a sufficiently large power of $q$. By Proposition 4.12, we see that $G^{i+1}$ satisfies Items 1 and 3, choosing $a^{i+1}_{j}=a^{i}_{j}$ for all $j\neq i+1$. We also choose $k$ to be large enough so that the orders of $a_{1}^{i},\ldots,a_{i}^{i}$ and their images $a_{1}^{i+1},\ldots,a_{i}^{i+1}$ have the same order, and each finite subgroup from $\mathcal{E}_{i}$ embeds into $G^{i+1}$. Moreover, by Proposition 4.12 again, this triple is Cohen–Lyndon, so Corollary 4.3 applies, and proves Item 2. Finally, let $E\coloneqq E_{G^{i}}(a_{i+1}^{i+1})/\langle(a_{i+1}^{i+1})^{k}\rangle$. Corollary 4.3 and property $(T)$ for $G^{i}$ give a short exact sequence
$0\to H_{2}(G^{i};V)\to H_{2}(G^{i+1};V)\to V_{E}\to 0,$
so setting $\mathcal{E}_{i+1}=\mathcal{E}_{i}\cup\{E\}$, we obtain the first part of Item 4. The second part follows again by induction, if we choose $k$ to be sufficiently large, and we conclude. ∎
@@
We proceed as in the proof of Theorem 6.1, using the prime $p\neq q$, but starting with the group $G_{0}$ from Lemma 6.2. The result will be a group $G_{x}$ with the same properties as in Theorem 6.1, with the exception that cohomological dimension now has to be taken modulo $\{p,q\}$, and $b^{(2)}_{2}(G_{x})$ is computed starting from
$b^{(2)}_{2}(G_{0}\to G_{x})=\dim_{\mathcal{N}(G_{x})}\Bigl(\bigoplus_{E\in\mathcal{E}}\mathcal{N}(G_{\alpha})_{E}\Bigr)=\sum\limits_{E\in\mathcal{E}}|E|^{-1}.$
We let this value be $\varepsilon<x$. The same computations go through, where now work with a base-$p$ representation of $x-\varepsilon$ instead. For the finiteness of the higher $\ell^{2}$-Betti numbers of $G_{x}$ we refer to Remark 6.3.
In each induction step in the proof of Theorem 6.1, we chose special elements $g_{i}\in G_{i}$, using Proposition 4.13 with arbitrary $x,H$. We will now explain how to use the full power of Proposition 4.13 in the construction so that this time the colimit is simple. Enumerate the set $\mathcal{A}\coloneqq\{a_{1},\ldots,a_{m}\}\times G_{0}$. When the group $G_{i}$ has been built, let $(a,h)$ denote the $i$-th element of $\mathcal{A}$. If $h$ is trivial, then we pick $g_{i}$ to be an arbitrary special element, and proceed as before. Otherwise, we use Proposition 4.13 to pick $g_{i}\in a^{-1}\left<\!\left<h\right>\!\right>_{G_{i}}$, which is possible since $\left<\!\left<h\right>\!\right>_{G_{i}}$ is suitable by Lemma 4.10. We claim that now $a\in\left<\!\left<h\right>\!\right>_{G_{i+1}}\leqslant G_{i+1}$. Indeed, consider the image of $a$ in the quotient $G_{i+1}/\left<\!\left<h\right>\!\right>_{G_{i+1}}$. On the one hand, it has order a power of $q$, because $a$ has order a power of $q$ in $G_{0}$. On the other hand, it equals the image of $g_{i}$, which has order a power of $p$ in $G_{i+1}$, by construction of the quotient $G_{i}\to G_{i+1}$. Because $p\neq q$, we see that $a$ must be trivial in $G_{i+1}/\left<\!\left<h\right>\!\right>_{G_{i+1}}$, which proves our claim.
As a result, for every $(a,h)\in\mathcal{A}$, either $h$ is trivial in $G_{x}$, or $a$ lies in the normal closure of $h$ in $G_{x}$. Because $a$ ranges over a generating set for $G_{0}$ and $h$ ranges over $G_{0}$, this shows that every non-trivial element of $G_{x}$ normally generates, that is, $G_{x}$ is simple. ∎
@@
Note that the constructions give some information about higher $\ell^{2}$-Betti numbers as well. Indeed, $G_{x}$ is obtained from a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, denoted $G$, by a sequence of group-theoretic Dehn fillings satisfying Corollary 4.3. Using this and Proposition 2.2 we have
$b^{(2)}_{n}(G_{x})=b^{(2)}_{n}(G\to G_{x})$
for all $n\geqslant 3$. Since $G$ is of type $F$, this implies that all $\ell^{2}$-Betti numbers of $G_{x}$ are finite. We know from Theorem 3.10 that $G$ has property $(T_{3})$, so Corollary 3.9 gives $b^{(2)}_{3}(G_{x})=0$.
Moreover, $G$ satisfies the Singer conjecture [Lüc02, Corollary 5.16 on p. 231], that is $b^{(2)}_{n}(G)=0$ for all $n\neq 8$ and $b^{(2)}_{8}(G)=\chi\geqslant 1$. This does not give information about $b^{(2)}_{n}(G\to G_{x})$, however, assuming Lück’s approximation conjecture [Lüc02, Chapter 13], and repeating the construction ensuring that $G\to G_{x}$ is injective on a large ball centred at the identity, we would have that $b^{(2)}_{n}(G\to G_{x})$ is close to $b^{(2)}_{n}(G)$. It would then follow that $b^{(2)}_{n}(G_{x})$ is close to 0, for $n\geqslant 4,n\neq 8$, and close to $\chi$ for $n=8$.
@@
Taking $V=\mathbb{C}$ to be the trivial unitary representation, the proofs of Theorems 6.1 and A, and Proposition 2.2, show that $H_{2}(G_{x};\mathbb{C})$ is infinite-dimensional. It follows that the groups $G_{\alpha}$ do not have type $FP_{2}(\mathbb{Q})$.
@@ 7 An elementary construction
We conclude with the elementary construction giving F. We start with two basic product formulas.
@@
Taking $\dim_{\mathcal{N}(G_{x})}$ and $n\leqslant 3$, the first and last term give $2b^{(2)}_{n}(R_{x}\times F_{2}^{3})=2b^{(2)}_{n-1}(R_{x}\times F_{2}^{3})=0$, by Proposition 3.11, so
@@
Let $(G_{x})_{x\in\mathbb{R}_{>0}}$ be as in F. Then the groups $G_{x}\ast\mathbb{Z}$ are pairwise non-measure equivalent.